# Competing color superconductivity and color Kondo effect in quark matter

### v1

@@ I Introduction
Quarks and gluons, as described by Quantum Chromodynamics (QCD) [1, 2, 3, 4] are typically strongly interacting and confined within protons and neutrons. They become weakly coupled (asymptotically free) under extremely high temperatures or densities, undergoing deconfinement and forming a Quark-Gluon Plasma (QGP). This occurs at temperatures above the pseudo-critical value $T_{c}\approx 155~{}\mathrm{MeV}$, as established by lattice QCD [5, 6] and confirmed in heavy-ion collisions at RHIC and the LHC [7, 8, 9, 10, 11]. At high baryon density ($\mu_{B}\gtrsim 500~{}\mathrm{MeV}$), deconfinement is instead expected to occur at significantly lower temperatures, in the range $T\sim 50\text{--}100~{}\mathrm{MeV}$, depending on the location of the critical point and the structure of the QCD phase diagram [12]. Below this range, QCD is expected to host a wealth of exotic, confined phases such as color superconductivity and quarkyonic matter [13, 14, 15], which are actively studied in the context of compact stars and dense matter experiments. The study of the phenomenologically relevant lower-density regions, where novel phases emerge, proved to be challenging due to strong coupling and inherent non-perturbative effects, making theoretical and experimental investigations particularly demanding [16, 17].
@@
One particularly compelling phase is color superconductivity [18, 19, 20]. In this phase, quarks near the Fermi surface form Cooper pairs through attractive interactions in color channels, leading to a finite energy gap in the excitation spectrum and a spontaneous breaking of color gauge symmetry [21]. Such a state is believed to play a crucial role in the inner cores of neutron stars [22, 23] and may be accessible in ultrarelativistic heavy-ion collisions [24], offering a window into the behavior of strongly interacting matter under extreme conditions. Color-superconducting quark matter has multiple phases because quarks, which come in three colors and six flavors (up, down, strange, charm, beauty, and top), can form Cooper pairs in various ways. Each pairing pattern breaks different symmetries, leading to distinct excitation spectra and transport properties. Predicting the favored pairing mechanism among the many possibilities is extremely challenging. In principle, a full QCD calculation in the strong coupling regime should determine the answer, but current lattice QCD methods have not yet achieved this. To gain quantitative insights, one often turns to the Nambu–Jona–Lasinio (NJL) model, a simplified version of QCD that captures key aspects of color superconductivity [25, 26]. However, since the model replaces dynamical gluons with an effective four-fermion interaction, the model is not renormalizable, and hence a UV cut-off $\Lambda$ is required from the outset such that this parameter define the energy scale below which the theory is valid [27].
@@
In parallel, recent theoretical studies have uncovered the possibility of a QCD analog of the Kondo effect [28, 29, 30]. In this scenario, heavy quarks such as charm or bottom act as impurities within a sea of light quarks (up, down, and strange) [31]. The non-Abelian gluon exchange between light and heavy quarks generates an interaction reminiscent of the traditional Kondo effect observed in condensed matter systems, where conduction electrons screen magnetic impurities [32, 33, 34, 35, 36]. In quark matter, this interaction leads to the formation of a Kondo gap and the emergence of mixed states, where light and heavy quarks combine into Bogoliubov quasiparticles [29, 37].
@@
where $c$ is a numerical constant determined by the pairing channel at the cut-off scale $\Lambda$.
In contrast, the presence of a heavy quark impurity allows for color exchange with the Fermi sea, resulting in logarithmic enhancement of the scattering amplitude at low energy—a hallmark of the QCD Kondo effect. The associated Kondo scale is given by
$\Lambda_{K}\sim\Lambda\exp\left(-\frac{\pi}{2N_{c}J}\right)\quad N_{c}=3,$ (2)
@@
To describe these phenomena in more detail we proceed to discuss the theory at the regime of high quark chemical potential $\mu\gg\Lambda_{\text{QCD}}$, where the QCD coupling becomes weak due to asymptotic freedom. In this regime, gluon-mediated interactions between quarks can be approximated by local four-fermion interactions, allowing an effective description in terms of an NJL-type model.
@@
$+g^{2}\int\mathrm{d}x^{4}\mathrm{d}y^{4}(\bar{\psi}_{i}\gamma^{\mu}\tau^{a}\psi_{i})(x)D_{\mu\nu}^{ab}(x,y;\epsilon)(\bar{\psi}_{j}\gamma_{\nu}\tau^{b}\psi_{j})(y)$
@@
$\mathcal{S}$ =${}_{\rm QCD-Kondo}=\mathcal{S}_{\rm q}$
$+g^{2}\int\mathrm{d}x^{4}\mathrm{d}y^{4}(\bar{\psi}_{i}\gamma^{0}\tau^{a}\psi_{i})(x)D_{00}^{ab}(x,y;\mu)T^{b}\delta(y),$ (5)
@@
In Ref. [40], the problem is studied in the limit $g=0$, where a S-wave approximation can be performed to write the effective theory in 1+1D as
$S^{2D}_{\text{eff}}=\int d^{2}x\left[\bar{\Psi}\left(i\Gamma^{\mu}\partial_{\mu}\right)\Psi+G\Psi\tau^{a}\Psi T^{a}\delta(x)\right],$ (6)
where the two dimensional Dirac matrices are given by $\Gamma^{0}=\sigma^{1}$, $\Gamma^{z}=-i\sigma^{2}$. $\Psi$ is a two-component quark field. In Ref. [40], this model Eq. (6) is studied using the conformal field theory technique. In this approach, the eigenstate of the bulk Hamiltonian with $U(3f)$ symmetry admits conformal embedding of the form [41]
@@
and the impurity only couples to the $SU(f)_{3}$ color degree of freedom, where, using the standard fusion rule, one can explicitly obtain physical quantities related to the impurities like impurity entropy, low-temperature susceptibility, and specific heat [42, 36].
In this work, we investigate the interplay between two competing phenomena—color superconductivity and the color Kondo effect—within a 1+1D toy model of QCD based on the SU(3) Gross-Neveu model [43, 44] with an impurity. Similar to the NJL model, the SU(3) Gross-Neveu model generates a superconducting mass gap through an attractive color density-density interaction. However, due to Coleman’s theorem, true spontaneous symmetry breaking is forbidden in 1+1D. Instead, the fermionic bilinear exhibits quasi-long-range order, with correlations that decay algebraically rather than exponentially, while the mass gap is generated dynamically. When a localized color impurity is placed in the system, a color Kondo effect takes place as discussed in terms of RG analysis. The rest of the paper firms up this assertion by providing a detailed exact solution of the model.
@@ II The Model
As a first step towards a controlled model, we shall consider a model with a single flavor of quark with attractive interaction between the quarks of opposite chirality, which dynamically opens a mass gap, and we shall consider a single heavy quark impurity localized at $x=0$ where, when the bulk scatters off the impurity, its chirality changes. We shall return to a more realistic description of the QCD Kondo problem in superconducting quark matter in a subsequent paper by considering a multiflavor problem with attractive color interaction between different intra- and inter-flavors. The Hamiltonian of the model under consideration is
$H$ $=-i\int_{-L}^{0}\Big{\{}\psi^{\dagger}_{+,a}(x)\partial_{x}\psi_{+,\alpha}(x)-\psi^{\dagger}_{-,a}(x)\partial_{x}\psi_{-,a}(x)\Big{\}}$
$+\int_{-L}^{0}\Big{\{}g\psi^{\dagger}_{+,a}(x)\vec{\tau}_{ab}\psi^{\dagger}_{+,b}(x)\psi^{\dagger}_{+,c}(x)\vec{\tau}_{cd}\psi^{\dagger}_{+,d}(x)$
@@
where $\psi_{\pm,a}$ is a fermionic annhilator operator where the first index $\pm$ denotes its chirality whereas the second index $a$ is the $SU(3)$ color index. The matrices $\tau$ are Gell-Mann matrices, $SU(3)$ generators that act on the color space of the quarks and $\vec{T}$ describes the localized impurity. The coupling constant $g$ is the attractive interaction between the bulk quarks of opposite chirality, whereas $J$ is the Kondo coupling.
We will describe the various phases of the Hamiltonian as functions of the couplings $J,g$. The bulk coupling constant $g$ flows to strong coupling where the model opens a superconducting gap [43, 44]. In the absence of $g$, the Kondo coupling $J$ also flows to strong coupling where the impurity is screened by a multiparticle cloud of bulk fermions [45, 34]. In this work, we shall provide an exact solution for the Hamiltonian, which allows the study of the interplay between the two phenomena when both couplings are present.
This interplay between superconductivity and the Kondo screening has long been a subject of study in condensed matter physics [46, 47, 48, 49, 50, 51, 52, 53, 54, 55, 56]. In a superconducting material, the fundamental state is characterized by the attractive interaction between electrons, leading to the formation of Cooper pairs and a globally coherent superconducting phase [57]. However, the introduction of a non-magnetic impurity can alter this picture locally. Instead of participating in the collective Cooper pairing, conduction electrons in the vicinity of the impurity become strongly correlated with it, forming a localized many-body singlet state analogous to the Kondo effect observed in normal metals containing magnetic impurities.
The competition between the bulk’s tendency to form Cooper pairs and the local tendency to screen the impurity through singlet formation gives rise to a fascinating array of behaviors in the vicinity of the impurity [58, 59, 60]. Depending on the relative strength of the coupling between the bulk superconducting electrons and the impurity, the impurity’s magnetic moment can be screened by a complex multi-particle cloud, characteristic of the Kondo effect. Or the impurity can induce localized bound states within the superconducting energy gap, known as Yu-Shiba-Rusinov (YSR) states, which effectively screen the impurity through single-particle excitations [46, 47, 48]. In yet other regimes, the impurity might remain entirely unscreened, retaining its local magnetic moment.
@@
Before presenting our main results, we note that recent advances in cold‐atom technology make it plausible to realize the toy model proposed here in the laboratory. In particular, the capability to engineer and probe local moments in multicomponent fermionic systems—thereby accessing a Kondo‐like coupling [61]—and the development of SU($N$)‐symmetric gases using alkaline-earth-like atoms have been demonstrated in Refs . [62, 63, 64]. Thus, it should be possible to implement our model within a suitably designed cold‐atom experiment in the near future.
@@ III Summary of main results
This shows that at an initial ultraviolet (UV) energy scale $D_{0}$, the behavior of the boundary coupling $J$ can be approximated as $\beta(J)\approx-\frac{6J}{\pi}(J-2g)$ and as discussed above it indeed comprises two competing effects: a standard Kondo-type term, $-CJ^{2}$ with $C=\frac{6}{\pi}$, which tends to drive $J$ towards strong coupling, and a bulk-induced term, $+agJ$ with $a=\frac{12}{\pi}$, which acts to suppress $J$. The fact that the bulk coupling $g$ flows to strong coupling as energy scales decrease, coupled with the boundary coupling $J$ exhibiting a sign and magnitude that are intricately tied to both $g$ and its own value, unveils a far more nuanced boundary physics than the standard Kondo effect in Fermi liquid theory. This interplay of bulk and boundary RG flow shapes the impurity’s fate, leading to a richer tapestry of potential behaviors.
@@
Figure 1: The weak-coupling RG flow diagram for Hamiltonian Eq. (8) are given by Eq. (9) and Eq. (10). The flow is away from the non-interacting point $J=0=g$. The magenta curves denote the Kondo phase, where the boundary coupling flows to strong coupling where the impurity is screened by a multiparticle Kondo cloud. The cyan lines indicate the YSR phase, where the impurity is screened by single-particle bound mode. Finally, the orange lines depict the unscreened regime, where the boundary coupling flows to zero such that the impurity is unscreened. The dashed black curves demarcate the three boundary phases of the model.
The renormalization group flow diagram depicted in Fig. 1 illustrates the evolution of the couplings $g$ and $J$ as the energy scale is lowered, originating from the non-interacting fixed point at $J=g=0$. These RG equations, derived using an unconventional cutoff scheme within the framework of the Bethe Ansatz solution and detailed in Appendix E, reveal distinct regions corresponding to different impurity behaviors. The flow lines shown in orange represent the parameter regimes where the boundary coupling $J$ flows towards weaker values in the infrared, ultimately leading to the impurity remaining as a local, unscreened moment. In contrast, the flow lines in magenta (or the color you perceive as indicative of strong coupling) correspond to the parametric regimes where the boundary coupling $J$ flows towards strong coupling, resulting in the screening of the impurity by a Kondo cloud of multiple particles. In particular, the flow lines in cyan illustrate the parametric regimes where the boundary coupling evolves towards the YRS phase characterized by screening the impurity through the formation of a single particle bound mode below the bulk mass gap.

### v2

@@ I Introduction
Quarks and gluons, as described by Quantum Chromodynamics (QCD) [1, 2, 3, 4] are typically strongly interacting and confined within protons and neutrons. They become weakly coupled (asymptotically free) under extremely high temperatures or densities, undergoing deconfinement and forming a Quark-Gluon Plasma (QGP). This occurs at temperatures above the pseudo-critical value $T_{c}\approx 155~\mathrm{MeV}$, as established by lattice QCD [5, 6] and confirmed in heavy-ion collisions at RHIC and the LHC [7, 8, 9, 10, 11]. At high baryon density ($\mu_{B}\gtrsim 500~\mathrm{MeV}$), deconfinement is instead expected to occur at significantly lower temperatures, in the range $T\sim 50\text{--}100~\mathrm{MeV}$, depending on the location of the critical point and the structure of the QCD phase diagram [12]. Below this range, QCD is expected to host a wealth of exotic, confined phases such as color superconductivity and quarkyonic matter [13, 14, 15], which are actively studied in the context of compact stars and dense matter experiments. The study of the phenomenologically relevant low-temperature regions, where novel phases emerge, proved to be challenging due to strong coupling and inherent non-perturbative effects, making theoretical and experimental investigations particularly demanding [16, 17].
@@
One particularly compelling phase is color superconductivity [18, 19, 20]. In this phase, quarks near the Fermi surface form Cooper pairs through attractive interactions in color channels, leading to a finite energy gap in the excitation spectrum and a spontaneous breaking of color gauge symmetry [21] as shown in the phase diagram 1. Such a state is believed to play a crucial role in the inner cores of neutron stars [22, 23] and may be accessible in ultrarelativistic heavy-ion collisions [24], offering a window into the behavior of strongly interacting matter under extreme conditions. Color-superconducting quark matter has multiple phases because quarks, which come in three colors and six flavors (up, down, strange, charm, beauty, and top), can form Cooper pairs in various ways. Each pairing pattern breaks different symmetries, leading to distinct excitation spectra and transport properties. Predicting the favored pairing mechanism among the many possibilities is extremely challenging. In principle, a full QCD calculation in the strong coupling regime should determine the answer, but current lattice QCD methods have not yet achieved this. To gain quantitative insights, one often turns to the Nambu–Jona–Lasinio (NJL) model, a simplified version of QCD that captures key aspects of color superconductivity [25, 26]. However, since the model replaces dynamical gluons with an effective four-fermion interaction, the model is not renormalizable, and hence a UV cut-off $\Lambda$ is required from the outset such that this parameter defines the energy scale below which the theory is valid [27].
Figure 1: Phase diagram of QCD in temperature and chemical potential phase showing the hadronic, quark-gluon plasma, and color superconducting phase.
@@
In parallel, recent theoretical studies have uncovered the possibility of a QCD analog of the Kondo effect [28, 29, 30, 31]. In this scenario, heavy quarks such as charm or bottom act as impurities within a sea of light quarks (up, down, and strange) [32]. The non-Abelian gluon exchange between light and heavy quarks generates an interaction reminiscent of the traditional Kondo effect observed in condensed matter systems, where conduction electrons screen magnetic impurities [33, 34, 35, 36, 37]. In quark matter, this interaction leads to the formation of a Kondo gap and the emergence of mixed states, where light and heavy quarks combine into Bogoliubov quasiparticles [29, 38].
@@
where $G$ is the 4-fermion attractive coupling, $c$ is a numerical constant determined by the pairing channel at the cut-off scale $\Lambda$.
In contrast, the presence of a heavy quark impurity allows for color exchange with the Fermi sea, resulting in logarithmic enhancement of the scattering amplitude at low energy—a hallmark of the QCD Kondo effect. The associated Kondo scale is related to cut-off scale $\Lambda$ by
$\Lambda_{K}\sim\Lambda\exp\left(-\frac{\pi}{2N_{c}J}\right),\quad N_{c}=3\,$, (2)
@@
To describe these phenomena in more detail, we proceed to discuss the theory in the regime of high quark chemical potential $\mu\gg\Lambda_{\text{QCD}}$, where the QCD coupling becomes weak due to asymptotic freedom. In this regime, gluon-mediated interactions between quarks can be approximated by local four-fermion interactions, allowing an effective description in terms of an NJL-type model.
@@
$+G\int\mathrm{d}x^{4}\mathrm{d}y^{4}(\bar{\psi}_{i}\gamma^{\mu}\tau^{a}\psi_{i})(x)D_{\mu\nu}^{ab}(x,y;\epsilon)(\bar{\psi}_{j}\gamma_{\nu}\tau^{b}\psi_{j})(y)$
@@
$\mathcal{S}$ $_{\rm QCD-Kondo}=\mathcal{S}_{\rm q}$
$+J\int\mathrm{d}x^{4}\mathrm{d}y^{4}(\bar{\psi}_{i}\gamma^{0}\tau^{a}\psi_{i})(x)D_{00}^{ab}(x,y;\mu)T^{b}\delta(y),$ (5)
@@
In Ref. [41], the problem is studied in the limit $g=0$, where an S-wave approximation can be performed to write the effective theory in 1+1D as
$S^{2D}_{\text{eff}}=\int d^{2}x\left[\bar{\Psi}\left(i\Gamma^{\mu}\partial_{\mu}\right)\Psi+J^{\prime}\bar{\Psi}\tau^{a}\Psi T^{a}\delta(x)\right],$ (6)
where the two-dimensional Dirac matrices are given by $\Gamma^{0}=\sigma^{1}$, $\Gamma^{z}=-i\sigma^{2}$ and $J^{\prime}$ is the effective Kondo coupling. $\Psi$ is a two-component quark field. In Ref. [41], this model Eq. (6) is studied using the conformal field theory technique. In this approach, the eigenstate of the bulk Hamiltonian with $U(3f)$ symmetry admits a conformal embedding of the form [42]
@@
and the impurity only couples to the $SU(3)_{f}$ color degree of freedom, where, using the standard fusion rule, one can explicitly obtain physical quantities related to the impurities like impurity entropy, low-temperature susceptibility, and specific heat [43, 37].
In this work, we investigate the interplay between two competing phenomena—color superconductivity and the color Kondo effect—within a 1+1D toy model of QCD based on the SU(3) Gross-Neveu model [44, 45] with an impurity. Similar to the NJL model, the SU(3) Gross-Neveu model generates a superconducting mass gap through an attractive color density-density interaction. However, due to Coleman’s theorem, true spontaneous symmetry breaking is forbidden in 1+1D. Instead, the fermionic bilinear exhibits quasi-long-range order, with correlations that decay algebraically rather than exponentially, while the mass gap is generated dynamically. When a localized color impurity is placed in the system, a color Kondo effect takes place as discussed in terms of RG analysis. The rest of the paper shows that, while the RG description is correct in the regime $\Lambda_{K}\gg\Delta$, the exact solution reveals that the model is significantly richer, exhibiting additional impurity phases.
@@ II The Model
As a first step towards a controlled model, we shall consider a model with a single flavor of quark with attractive interaction between the quarks of opposite chirality, which dynamically opens a mass gap, and we shall consider a single heavy quark impurity localized at $x=0$ where, when the bulk scatters off the impurity, its chirality changes. This is pictorially shown in Fig. 2. We shall return to a more realistic description of the QCD Kondo problem in superconducting quark matter in a subsequent paper by considering a multiflavor problem with attractive color interaction between different intra- and inter-flavors. The Hamiltonian of the model under consideration is
$H$ $=-i\int_{-L}^{0}\mathrm{d}x\Big{\{}\psi^{\dagger}_{+,a}(x)\partial_{x}\psi_{+,a}(x)-\psi^{\dagger}_{-,a}(x)\partial_{x}\psi_{-,a}(x)\Big{\}}$
$+\int_{-L}^{0}\mathrm{d}x\Big{\{}g\,\psi^{\dagger}_{+,a}(x)\vec{\tau}_{ab}\psi_{+,b}(x)\,\psi^{\dagger}_{-,c}(x)\vec{\tau}_{cd}\psi_{-,d}(x)$
@@
where $\psi_{\pm,a}$ is a fermionic annihilator operator where the first index $\pm$ denotes its chirality whereas the second index $a$ is the $SU(3)$ color index. The matrices $\tau$ are Gell-Mann matrices, $SU(3)$ generators that act on the color space of the quarks, and $\vec{T}$ describes the localized impurity. The coupling constant $g$ is the attractive interaction between the bulk quarks of opposite chirality, whereas $J$ is the Kondo coupling.
We will describe the various phases of the Hamiltonian as functions of the couplings $J$ and $g$. The bulk coupling constant $g$ flows to strong coupling where the model opens a superconducting gap [44, 45]. In the absence of $g$, the Kondo coupling $J$ also flows to strong coupling where the impurity is screened by a multiparticle cloud of bulk fermions [46, 35]. In this work, we shall provide an exact solution for the Hamiltonian, which allows the study of the interplay between the two phenomena when both couplings are present.
This interplay between superconductivity and the Kondo screening has long been a subject of study in condensed matter physics [47, 48, 49, 50, 51, 52, 53, 54, 55, 56, 57]. In a superconducting material, the fundamental state is characterized by the attractive interaction between electrons, leading to the formation of Cooper pairs and a globally coherent superconducting phase [58]. However, the introduction of a magnetic impurity can alter this picture locally. Instead of participating in the collective Cooper pairing, conduction electrons in the vicinity of the impurity become strongly correlated with it, forming a localized many-body singlet state analogous to the Kondo effect observed in normal metals containing magnetic impurities.
The competition between the bulk’s tendency to form Cooper pairs and the local tendency to screen the impurity through singlet formation gives rise to a fascinating array of phenomena in the vicinity of the impurity [59, 60, 61]. Depending on the relative strength of the coupling between the bulk superconducting electrons $g$ and the impurity $J$, the impurity’s magnetic moment can be screened by a complex multi-particle cloud, characteristic of the Kondo effect. Or the impurity can induce localized bound states within the superconducting energy gap, known as Yu-Shiba-Rusinov (YSR) states, which effectively screen the impurity through single-particle excitations [47, 48, 49]. In yet other regimes, the impurity might remain entirely unscreened, retaining its local magnetic moment.
@@
[…]
Figure 2: Cartoon depiction of the model Eq. (8) where left and right moving chiral fermions with marginal bulk attracting current-current interaction $g$ and open boundary conditions are coupled to an SU(3) moment at the boundary with Kondo coupling $J$.
Before presenting our main results, we note that recent advances in cold atom technology make it plausible to realize the toy model proposed here in the laboratory. In particular, the capability to engineer and probe local moments in multicomponent fermionic systems—thereby accessing a Kondo‐like coupling [62]—and the development of SU(NN)‐symmetric gases using alkaline-earth-like atoms have been demonstrated in Refs. [63, 64, 65]. Thus, it should be possible to implement our model within a suitably designed cold atom experiment in the near future by realizing the setup shown in schematic Fig.2.
@@ III Summary of main results
This shows that at an initial ultraviolet (UV) energy scale $D_{0}$, the behavior of the boundary coupling $J$ can be approximated as $\beta(J)\approx-\frac{6J}{\pi}(J-2g)$ and as discussed above it indeed comprises two competing effects: a standard Kondo-type term, $CJ^{2}$ with $C=\frac{6}{\pi}$, which tends to drive $J$ towards strong coupling, and a bulk-induced term, +agJ+agJ with $a=\frac{12}{\pi}$, which acts to suppress $J$. The fact that the bulk coupling $g$ flows to strong coupling as energy scales decrease, coupled with the boundary coupling $J$ exhibiting a sign and magnitude that are intricately tied to both $g$ and its own value, unveils a far more nuanced boundary physics than the standard Kondo effect in Fermi liquid theory. This interplay of bulk and boundary RG flow shapes the impurity’s fate, leading to a richer tapestry of potential behaviors.
@@
Figure 3: The weak-coupling RG flow diagram for Hamiltonian Eq. (8) are given by Eq. (9) and Eq. (10). The flow is away from the non-interacting point $J=0=g$. The magenta curves denote the Kondo phase, where the boundary coupling flows to strong coupling, and the impurity is screened by a multiparticle Kondo cloud. The cyan lines indicate the YSR phase, where the impurity is screened by a single-particle bound mode. Finally, the orange lines depict the unscreened regime, where the boundary coupling flows to zero, leaving the impurity unscreened. The dashed black curves demarcate the three boundary phases of the model.
The renormalization group flow diagram depicted in Fig. 3 illustrates the evolution of the couplings $g$ and $J$ as the energy scale is lowered, originating from the non-interacting fixed point at $J=g=0$. These RG equations, derived using an unconventional cutoff scheme within the framework of the Bethe Ansatz solution and detailed in Appendix E, reveal distinct regions corresponding to different impurity behaviors. The flow lines shown in orange represent the parameter regimes where the boundary coupling $J$ flows towards weaker values in the infrared, ultimately leading to the impurity remaining as a local, unscreened moment. In contrast, the flow lines in magenta (or the color you perceive as indicative of strong coupling) correspond to the parametric regimes where the boundary coupling $J$ flows towards strong coupling, resulting in the screening of the impurity by a Kondo cloud of multiple particles. In particular, the flow lines in cyan illustrate the parametric regimes where the boundary coupling evolves towards the YSR phase characterized by screening the impurity through the formation of a single particle bound mode below the bulk mass gap.